The length of a rectangle is 3in. more than twice its width. If the perimeter of the rectangle is 36in., find the dimensions of the rectangle.
Answer by checkley71(8403)

PERIMETER =2*LENGTH+2*WIDTH OR LENGTH IS 2W+3 AND W=WIDTH THUS
2(2W+3)+2W=36 OR 4W+6+2W=36 OR 6W=30 OR W=30/6 OR W=5
PROOF 2*5+2(2*5+3)=36 OR 10+2(13)=36 OR 10+26=36 OR 36=36